A monopolist faces two totally separated markets with inverse demand p=100 – qA and
p=160−2qB respectively. The monopolist has no fixed costs and a marginal cost given by mc= 2 /3
q Find the profit maximizing total output and how much of it that is sold on market A and market
B respectively if the monopoly uses third degree price discrimination.
1
Expert's answer
2020-06-11T10:56:15-0400

Third degree price discrimination involves monopolists having the ability to sell commodities at different prices in different markets.

Market A

P=100 – QA

Total Revenue (TR) = "P\times Q"

TRA = "(100 – Q_{A}) \times Q_{A}"

TRA = "100 Q_{A} \u2013 Q_{A}^2"

MRA ="\varDelta TR_{A} / \varDelta Q_{A} = \varDelta(100Q_{A} – Q_{A}^2)/ \varDelta Q_{A}"

MRA = 100 – 2QA

Market B

P=160−2QB

Total Revenue (TR) = "P \times Q"

TRB = "(160−2Q_{b}) \times Q_{b}"

TRB = "160Q_{b} \u22122Q_{b}^2"

MRB ="\varDelta TR_{b}/ \varDelta Q_{b} = \varDelta (160Q_{b} −2Q_{b}^2)/ \varDelta Q_{b}"

MRB = 160 – 4QB


Equilibrium: MR = MC

Market A

MRA = 100 – 2QA = "\frac 23 Q_{A}"

"\frac 23" QA + 2QA = 100

"\frac 83Q_{A}" = 100

QA = "100 \times \frac 38"

QA = 37.5

P=100 – QA

P=100 – 37.5

P= 62.5


Total Revenue

TRA = "100Q_{A} \u2013 Q_{A}^2"

TRA = 100(37.5) – 37.52

TRA = 2,343.75


Market B

MRB = 160 – 4QB ="\frac 23Q_{b}"

"\frac {14}{3}" QB = 100

Q = "100 \times \frac {3}{14}"

QB = 21.43


Price

P="160\u22122Q_{b}"

P= "160 − 2 \times 21.43"

P = 160 – 42.86

P = 117.14


Total Revenue

TRB = "160Q_{b} \u22122Q_{b}^2"

TRB = 160(21.43) – 2(21.432)

TRB = 3,428.8 – 918.4898

TRB = 2,510.31